What is the prime factorization of 72?
Answer by jim_thompson5910(21685)

Solved by pluggable solver: Factor any number
72 is NOT a prime number: 72 = 2 * 2 * 2 * 3 * 3

Work Shown

72 is divisible by 2: 72 = 36 * 2.
36 is divisible by 2: 36 = 18 * 2.
18 is divisible by 2: 18 = 9 * 2.
9 is divisible by 3: 9 = 3 * 3.
3 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.